# GyroSwin: 5D Surrogates for Gyrokinetic Plasma Turbulence Simulations

### v1

@@ 1 Introduction
Nuclear fusion promises sustainable energy by fusing hydrogen isotopes in an ionized plasma. The plasma must be confined in a magnetic field as it reaches hundreds of millions of degrees. During this process, turbulence usually arises due to microinstabilities, resulting in energy and particle transport towards the reactor walls. The design and control of plasma scenarios strictly require knowledge of turbulent transport, which can be obtained via nonlinear gyrokinetic simulations that evolve a 5D Partial Differential Equation (PDE) over time.
@@ 2 Background and Related Work
Plasma kinetics models the time evolution of electrons and ions via the distribution function $f$ based on 3D coordinates, their parallel and perpendicular velocities, and the angle around the field lines. Gyrokinetics (Krommes, 2012) is a reduced form of plasma kinetics that is more efficient by averaging out the angle around the field lines and only considering the guiding center of a particle. Local gyrokinetics focuses on perpendicular spatial scales comparable to the gyroradius and on frequencies much lower than the particle cyclotron frequency, the circular motion of charged particles around magnetic field lines due to the Lorentz force. Hence, the 5D gyrokinetic distribution function can be written as $f=f(k_{x},k_{y},s,v_{||},\mu)$, where $k_{x}$ and $k_{y}$ are spectral coordinates (for the spatial $x$ and $y$), $s$ is the toroidal coordinate along a field line, and $v_{||}$ and $\mu$ represent parallel and perpendicular velocity components, respectively. The perturbed time-evolution of $f$, for each species (ions and electrons), is governed by
@@
$Q=\int\mathbf{C}\int\mathbf{v}^{2}\bm{\phi}\bm{f}\>\mathrm{d}v_{\parallel}\mathrm{d}\mu\>\>\mathrm{d}x\mathrm{d}y\mathrm{d}s,\bm{\phi}=\mathbf{A}\int\mathbf{J_{0}}\bm{f}\>\mathrm{d}v_{\parallel}\mathrm{d}\mu,\quad$ (2)
@@ 6 Conclusions and Limitations
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We acknowledge EuroHPC Joint Undertaking for awarding us access to Leonardo at CINECA, Italy, and Deucalion at MACC, Portugal.
@@ Appendix B Quasilinear models
To fit the nomalisation factor $C$, we assume access to a vector of nonlinear fluxes $\bm{q}\in\mathbb{R}^{j}$ and set the right-hand side of Equation˜20 to $\bm{x}\in\mathbb{R}^{j}$, then the optimal solution for $C$ can be obtained via a least-squares fit
@@ Appendix C Data Generation and Visualisation
In this work we mainly consider turbulence driven by ion temperature gradients. We leverage the numerical code GKW (Peeters et al., 2009) for generating nonlinear and linear Gyrokinetic simulations, by varying four parameters: $R/L_{t}$, $R/L_{n}$, $\hat{s}$, and $q$, which significantly affect emerging turbulence in the Plasma. $R/L_{t}$ is the ion temperature gradient, which is the main driver of turbulence. Larger values of $R/L_{t}$ always result in stronger turbulence (see Figure˜5). $R/L_{n}$ is the density gradient, whose effect is less pronounced. It can have a stabilizing effect, but can sometimes also lead to enhanced turbulence. The parameter $q$ denotes the so-called safety factor, which is the inverse of the rotational transform and describes how often a particle takes a poloidal turn before taking a toroidal turn. $\hat{s}$ denotes magnetic shearing, hence it usually has a stabilizing effect as more magnetic shearing results in improved isolation of the Plasma. We specify the ranges for sampling the four parameters as $R/L_{T}\in[1,12]$, $R/L_{n}\in[1,7]$, $q\in[1,9]$, and $\hat{s}\in[0.5,5]$. Additionally, we also vary the noise amplitude of the initial condition within [1e5,1e3][1e-5,1e-3] and the initial condition itself as one of $\operatorname{sin},\operatorname{cos},\operatorname{random}]$. We first generated the nonlinear simulations and then use the same parameter combination for each linear simulation. Thel linear simulations are required for implementing the quasilinear model.
@@ D.1 Baselines
FNO. The neuraloperator library [^3] https://github.com/neuraloperator/neuraloperator (Kovachki et al., 2023) contains tensorized nD implementations of Fourier convolution layers. However, as shown in LABEL:app:performance these are too expensive to be reliably trained on our data. To enable training FNO (Li et al., 2020), we utilize a much faster 3D FNO which operates only on the spatial dimensions, while flattening the velocity space in the channels. This baseline has 256 latent dimension, 4 layers and considers $\frac{1}{2}$ of the total modes on each dimension (5,004 flat spatial modes).

### v2

@@ 1 Introduction
Nuclear fusion promises sustainable energy by fusing hydrogen isotopes within an ionized plasma. Because the plasma reaches temperatures of hundreds of millions of degrees, it must be confined by a magnetic field. During confinement, turbulence can arise from micro-instabilities, leading to energy and particle transport toward the reactor walls. This causes plasma to leak from its magnetic cage, resulting in heat and density loss, and therefore impaired energy production. The design and control of plasma scenarios strictly require knowledge of turbulent transport, which can be obtained via nonlinear gyrokinetic simulations that evolve a 5D Partial Differential Equation (PDE) over time.
@@ 2 Background and Related Work
Plasma kinetics models the time evolution of electrons and ions via the distribution function $f$ based on 3D coordinates, their parallel and perpendicular velocities, and the angle around the field lines. Gyrokinetics (Krommes, 2012) is a reduced form of plasma kinetics that is more efficient by averaging out the angle around the field lines and only considering a particle’s guiding center. Local gyrokinetics focuses on perpendicular spatial scales comparable to the gyroradius and on frequencies much lower than the particle cyclotron frequency, the circular motion of charged particles around magnetic field lines due to the Lorentz force. Hence, the 5D gyrokinetic distribution function can be written as $f=f(k_{x},k_{y},s,v_{||},\mu)$, where $k_{x}$ and $k_{y}$ are spectral coordinates (for the spatial $x$ and $y$), $s$ is the toroidal coordinate along a field line, and $v_{||}$ and $\mu$ represent parallel and perpendicular velocity components, respectively. The perturbed time-evolution of $f$, for each species (ions and electrons), is governed by
@@
$Q=\int\mathbf{C}\int\mathbf{v}^{2}\bm{\phi}\bm{f}\>\mathrm{d}v_{\parallel}\mathrm{d}\mu\>\>\mathrm{d}x\mathrm{d}y\mathrm{d}s,\qquad\bm{\phi}=\mathbf{A}\int\mathbf{J_{0}}\bm{f}\>\mathrm{d}v_{\parallel}\mathrm{d}\mu,\quad$ (2)
@@ 6 Conclusions and Limitations
This work has been funded by the Fusion Futures Programme. As announced by the UK Government in October 2023, Fusion Futures aims to provide holistic support for the development of the fusion sector. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We acknowledge EuroHPC Joint Undertaking for awarding us access to Leonardo at CINECA, Italy, and Deucalion at MACC, Portugal.
@@ Appendix B Quasilinear models
To fit the normalisation factor $C$, we assume access to a vector of nonlinear fluxes $\bm{q}\in\mathbb{R}^{j}$ and set the right-hand side of Equation˜20 to $\bm{x}\in\mathbb{R}^{j}$, then the optimal solution for $C$ can be obtained via a least-squares fit
@@ Appendix C Data Generation and Visualisation
In this work, boundary conditions okay we mainly consider turbulence driven by ion temperature gradients. We leverage the numerical code GKW (Peeters et al., 2009) for generating nonlinear and linear Gyrokinetic simulations, by varying four parameters: $R/L_{t}$, $R/L_{n}$, $\hat{s}$, and $q$, which significantly affect emerging turbulence in the Plasma. $R/L_{t}$ is the ion temperature gradient, which is the main driver of turbulence. Larger values of $R/L_{t}$ always result in stronger turbulence (see Figure˜5). $R/L_{n}$ is the density gradient, whose effect is less pronounced. It can have a stabilizing effect, but can sometimes also lead to enhanced turbulence. The parameter $q$ denotes the so-called safety factor, which is the inverse of the rotational transform and describes how often a particle takes a poloidal turn before taking a toroidal turn. $\hat{s}$ denotes magnetic shearing, hence it usually has a stabilizing effect as more magnetic shearing results in improved isolation of the Plasma. We specify the ranges for sampling the four parameters as $R/L_{T}\in[1,12]$, $R/L_{n}\in[1,7]$, $q\in[1,9]$, and $\hat{s}\in[0.5,5]$. Additionally, we also vary the noise amplitude of the initial condition within [1e5,1e3][1e-5,1e-3] and the initial condition itself as one of $\operatorname{sin},\operatorname{cos},\operatorname{random}]$. We first generated the nonlinear simulations and then use the same parameter combination for each linear simulation. Thel linear simulations are required for implementing the quasilinear model.
@@ D.1 Baselines
FNO. The neuraloperator library [^3] https://github.com/neuraloperator/neuraloperator (Kovachki et al., 2023) contains tensorized nD implementations of Fourier convolution layers. However, as shown in Table˜3 these are too expensive to be reliably trained on our data. To enable training FNO (Li et al., 2020), we utilize a much faster 3D FNO which operates only on the spatial dimensions, while flattening the velocity space in the channels. This baseline has 256 latent dimension, 4 layers and considers $\frac{1}{2}$ of the total modes on each dimension (5,004 flat spatial modes).